### v4

@@ 3.5 BGLSS versus proposed model
Figure 12: Partial functional coefficients ($3^{\text{rd}}$ and $5^{\text{th}}$, true in red and estimated for the replications in which the coefficients were selected in black) from the BGLSS method, according to the data dispersion degree.
@@ 4 Application to real data and the GCV criterion
We considered two databases to evaluate the performance of the proposed model on a set of real data: the first with COVID-19 data freely available through the Brasil.IO accessible open data portal ([18]) for the definition of the response variable of the study; and the second with the socioeconomic data of the states and the Federal District made available by the IBGE (Brazilian Institute of Geography and Statistics) for the definition of the covariates for our analysis.
@@
Next, nine covariates were defined ($p=9$) as explanatory variables, which are:
  • demographic density;
  • HDI (Human Development Index);
  • Gini index of the per capita household income distribution;
  • unemployment rate of people aged 14 or over;
  • QoLLI (Quality of Life Loss Index);
  • percentage of population with health insurance;
  • number of hospital beds per 10 000 inhabitants;
  • number of doctors per 10 000 inhabitants;
  • average real income from the main job usually earned per month by persons aged 14 years or over (income).
The design matrix was defined with the covariates positioned in the same order in which they were presented above, so we have $Z_{1}\beta_{1}(.)$ as the functional coefficient associated with the demographic density, $Z_{2}\beta_{2}(.)$ as a functional coefficient associated with the HDI, and so on. In addition, the covariates were standardized by their mean and standard deviation before applying the model (see Section 1 of the Supplementary Material). In addition, the intercept was estimated as described in Section 3.1.
The Gibbs sampler implemented the proposed Bayesian model using two chains to enable a convergence diagnosis. The initialization of the chains, the maximum number of iterations, the choice of the burn-in period, the spacing of points between sampled values, as well as the choice of the hyperparameters and obtaining the Bayesian estimates were all established following what was described in Section 3.1 for the studies with synthetic data. Furthermore, we considered the model with $\mu$ as a parameter in this study.
Before presenting any results, it is worth mentioning that B-splines were used for the basis expansion of all partial functional coefficients, considering three values for $K$ ($K=5$, $K=10$ and $K=15$). The diagnostic analysis based on the method proposed by [12] attested for the convergence of the chains of the partial coefficients in all tested model configurations after the period of burn-in.
Unfortunately, it is impossible to calculate the MSE between the points of the estimated curve and those of the true curve in a study with real data since the true curve is unknown. Since calculating the MSE from the distances between the observed values and the points on the estimated curve is not recommended due to the possibility of overfitting, we consider an alternative measure based on a cross-validation procedure. Although the conventional cross-validation associated with the MSE is an effective procedure to overcome the problem of overfitting, it often becomes inefficient due to the high computational cost. As an alternative to the cross-validation procedure, there is the Generalized Cross Validation (GCV) ([14]). Adapting the GCV criterion presented by [14] to the reality of the proposed model, we obtain
@@
where $\tilde{y}_{ij}=y_{ij}-\tilde{\beta}_{0}(t_{ij})$, $\boldsymbol{S_{K}}$ is the projection matrix of the proposed model with $K$ bases on the expansions of the functional coefficients, $\boldsymbol{\hat{O}_{...}}$ is the estimate for $\boldsymbol{O_{...}}$ (as defined in Section 2.1) and $\boldsymbol{\hat{b}}$ is the estimate for $\boldsymbol{b}$.
@@
as shown in the details of the full conditional distribution for $\boldsymbol{b}$ in Section 2.1. Finally, when considering $\boldsymbol{\hat{b}}=\boldsymbol{\hat{Q}}^{-1}\boldsymbol{\hat{O}_{...}}^{{}^{\prime}}\boldsymbol{\tilde{y}_{..}}$, in which $\boldsymbol{\hat{Q}}$ and $\boldsymbol{\hat{O}_{...}}$ are the respective estimates of $\boldsymbol{Q}$ (also defined in Section 2.1) and $\boldsymbol{O_{...}}$, then the estimated curve is represented by
@@
Table 8 presents the value of the metric in (15) and the GCV criterion for each $K$ number of basis functions considered. We can observe that the fit quality is better with $K=5$, regardless of the metric used for analysis. Because the functional response data are observational, it is natural that the metric in (15) is not so close to one as they were in the simulation studies. Still, even with a reasonably low number of curves ($m=27$), the model with $K=5$ can return a value of almost 45% for that metric.
@@
Observing the posterior samples of each $Z_{l}$ (not shown here), we notice that there was perfect agreement between the points of the posterior samples and the summary statistics since the respective samples for $\hat{Z}_{l}$’s equal to zero are entirely characterized by zeros. Similarly, the respective samples for $\hat{Z}_{l}$’s equal to one are entirely characterized by ones. This reinforces the little uncertainty that the model has to define which variables should remain and which ones should be excluded from the model.
The Gini index and QoLLI covariates were the most influential for predicting the response variable since their respective functional coefficients are found on a larger scale. It is worth remembering that any analysis of the results must take into account that they are conditioned to some initial determining factors, such as the initial set of candidate covariates in the model and the hypothesis of linearity.
@@ 5 Final considerations
This work focuses on functional data modeling and proposes a model-based approach under the Bayesian paradigm for variable selection in FOSR. The proposed methodology shows low sensitivity to the hyperparameter $\mu$, which is desirable since this component does not control the regularization. On the other hand, $\lambda$ plays the role of a regularization hyperparameter, and naturally, the model has some sensitivity to it. If desired, it is still possible to add a Gamma prior to $\lambda^{2}$, leading to an improper prior and guaranteeing minor sensitivity of the model to $\lambda$. However, in this case, $\lambda$ would no longer play the role of regularizer and uniqueness problems in the stationary distribution could appear if there is multicollinearity between the covariates.
Through simulations, the proposed model proved superior in correctly selecting the covariates compared with other variable selection methods in FOSR. At the same time, it maintained an optimal level of goodness of fit. The competing methods in the simulation scenario with the highest data dispersion ($\sigma=20$) could not balance selection accuracy well with the goodness of fit. Therefore, they need to lose goodness of fit and increase regularization to ensure better accuracy in the selection.
@@ Acknowledgments
  • The real data sets are available at the Brasil.IO accessible open data portal ([18]) and IBGE data portal;

### v5

@@ 3.3.1 Comparative analysis with grLASSO, grSCAD, grMCP and BGLSS
Figure 5: Partial functional coefficients ($3^{\text{rd}}$ and $5^{\text{th}}$, true in red and estimated for the replications in which the coefficients were selected in black), according to the model used, considering data with $\sigma=0.2$.
(a) $\hat{\beta}_{3}(.)$ (proposed model, $\mu=0.5$).
(b) $\hat{\beta}_{5}(.)$ (proposed model, $\mu=0.5$).
(c) $\hat{\beta}_{3}(.)$ (proposed model, $\mu$ as a parameter).
(d) $\hat{\beta}_{5}(.)$ (proposed model, $\mu$ as a parameter).
(e) $\hat{\beta}_{3}(.)$ (grLASSO, $\lambda=0.001$).
(f) $\hat{\beta}_{5}(.)$ (grLASSO, $\lambda=0.001$).
(g) $\hat{\beta}_{3}(.)$ (grMCP, $\lambda=0.500$).
(h) $\hat{\beta}_{5}(.)$ (grMCP, $\lambda=0.500$).
(i) $\hat{\beta}_{3}(.)$ (grSCAD, $\lambda=0.001$).
(j) $\hat{\beta}_{5}(.)$ (grSCAD, $\lambda=0.001$).
(k) $\hat{\beta}_{3}(.)$ (BGLSS).
(l) $\hat{\beta}_{5}(.)$ (BGLSS).
Figure 6: Partial functional coefficients ($3^{\text{rd}}$ and $5^{\text{th}}$, true in red and estimated for the replications in which the coefficients were selected in black), according to the model used, considering data with $\sigma=20$.
@@ 4 Application to real data and the GCV criterion
We considered two databases to evaluate the performance of the proposed model on a set of real data: the first with COVID-19 data freely available through the Brasil.IO accessible open data portal (Justen and others, 2021) for the definition of the response variable of the study; and the second with the socioeconomic data of the states and the Federal District made available by the IBGE (Brazilian Institute of Geography and Statistics) for the definition of the covariates for our analysis.
@@
Next, nine covariates were defined ($p=9$) as explanatory variables, which are demographic density; HDI (Human Development Index); Gini index of the per capita household income distribution; unemployment rate of people aged 14 or over; QoLLI (Quality of Life Loss Index); percentage of population with health insurance; number of hospital beds per 10 000 inhabitants; number of doctors per 10 000 inhabitants; average real income from the main job usually earned per month by persons aged 14 years or over (income).
The design matrix was defined with the covariates positioned in the same order in which they were presented above, so we have $Z_{1}\beta_{1}(.)$ as the functional coefficient associated with the demographic density, $Z_{2}\beta_{2}(.)$ as a functional coefficient associated with the HDI, and so on. In addition, the covariates were standardized by their mean and standard deviation before applying the model (see Section 1 of the Supplementary Material). In addition, the intercept was estimated as described in Subsection 3.1.
The Gibbs sampler implemented the proposed Bayesian model using two chains to enable a convergence diagnosis. The initialization of the chains, the maximum number of iterations, the choice of the burn-in period, the spacing of points between sampled values, as well as the choice of the hyperparameters and obtaining the Bayesian estimates were all established following what was described in Subsection 3.1 for the studies with synthetic data. Furthermore, we considered the model with $\mu$ as a parameter in this study.
Before presenting any results, it is worth mentioning that cubic B-splines were used for the basis expansion of all partial functional coefficients, considering three values for $K$ ($K=5$, $K=10$ and $K=15$). The diagnostic analysis based on the method proposed by Gelman and Rubin (1992) indicated convergence of the chains of partial coefficients across all tested model configurations after the burn-in period.
Unfortunately, it is impossible to calculate the MSE between the points of the estimated curve and those of the true curve in a study with real data since the true curve is unknown. Since calculating the MSE from the distances between the observed values and the points on the estimated curve is not recommended due to the possibility of overfitting, we consider an alternative measure based on a cross-validation procedure. Although the conventional cross-validation associated with the MSE is an effective procedure to overcome the problem of overfitting, it often becomes inefficient due to the high computational cost. As an alternative to the cross-validation procedure, there is the Generalized Cross Validation (GCV) (Golub et al., 1979). Adapting the GCV criterion presented by Golub et al. (1979) to the reality of the proposed model, we obtain
@@
where $\tilde{y}_{ij}=y_{ij}-\tilde{\beta}_{0}(t_{ij})$, $\boldsymbol{S_{K}}$ is the projection matrix of the proposed model with $K$ bases on the expansions of the functional coefficients, $\boldsymbol{\hat{O}_{...}}$ is the estimate for $\boldsymbol{O_{...}}$ (as defined in Subsection A) and $\boldsymbol{\hat{b}}$ is the estimate for $\boldsymbol{b}$.
@@
as shown in the details of the full conditional distribution for $\boldsymbol{b}$ in Subsection A. Finally, when considering $\boldsymbol{\hat{b}}=\boldsymbol{\hat{Q}}^{-1}\boldsymbol{\hat{O}_{...}}^{{}^{\prime}}\boldsymbol{\tilde{y}_{..}}$, in which $\boldsymbol{\hat{Q}}$ and $\boldsymbol{\hat{O}_{...}}$ are the respective estimates of $\boldsymbol{Q}$ (also defined in Subsection A) and $\boldsymbol{O_{...}}$, then the estimated curve is represented by
@@
Table 4 presents the value of the metric in (7) and the GCV criterion for each $K$ number of basis functions considered. We can observe that the fit quality is better with $K=5$, regardless of the metric used for analysis. Because the functional response data are observational, it is natural that the metric in (7) is not so close to one as it was in the simulation studies. Still, even with a relatively low number of curves ($m=27$), the model with $K=5$ can achieve almost 45% on that metric.
@@
Observing the posterior samples of each $Z_{l}$ (not shown here), we notice that there was perfect agreement between the points of the posterior samples and the summary statistics since the respective samples for $\hat{Z}_{l}$’s equal to zero are entirely characterized by zeros. Similarly, the respective samples for $\hat{Z}_{l}$’s equal to one are entirely characterized by ones. This reinforces the model’s limited uncertainty about which variables to retain and which to exclude.
The Gini index and QoLLI covariates were the most influential in predicting the response variable, as their respective functional coefficients are larger in magnitude. It is worth remembering that any analysis of the results must take into account that they are conditioned to some initial determining factors, such as the initial set of candidate covariates in the model and the hypothesis of linearity.
@@ 5 Final considerations
This work focuses on functional data modelling and proposes a model-based approach under the Bayesian paradigm for variable selection in FOSR. The proposed methodology shows low sensitivity to the hyperparameter $\mu$, which is desirable since this component does not control the regularization. On the other hand, $\lambda$ plays the role of a regularization hyperparameter, and naturally, the model has some sensitivity to it. If desired, it is still possible to add a Gamma prior to $\lambda^{2}$, leading to an improper prior and guaranteeing minor sensitivity of the model to $\lambda$. However, in this case, $\lambda$ would no longer play the role of regularizer and uniqueness problems in the stationary distribution could appear if there is multicollinearity between the covariates.
Through simulations, the proposed model proved superior in correctly selecting the covariates compared with other variable selection methods in FOSR. At the same time, it maintained an optimal level of goodness-of-fit. The frequentist competing methods in the simulation scenario with the highest data dispersion ($\sigma=20$) could not balance selection accuracy well with the goodness of fit. Therefore, they need to sacrifice goodness of fit and increase regularization to achieve better accuracy in selection.
@@
We conducted a brief simulation study to compare the computational costs of our proposed model with its Bayesian competitor, BGLSS, which was also the most competitive alternative overall. For this comparison, the Gibbs sampler outlined in Panel 2.1 was implemented both with the nimble package (de Valpine et al., 2024) in R and through a naive implementation, which is more pedagogical in nature but computationally less efficient than the nimble-based version. Ten simulated datasets were generated according to Section 3.1 with $m=10$, $n=25$, $p=6$, $K=10$. Running posterior inference for all ten replications required 29 minutes for the nimble-based implementation of the proposed model, whereas the naive implementation took 14 hours and 40 minutes. In contrast, BGLSS took 2 minutes and 55 seconds. All runs were performed on a computer equipped with an 11th Gen Intel(R) Core(TM) i7-1195G7 @ 2.90GHz processor, using R version 4.4.1. Given that the nimble-based implementation took 29 minutes for 10 datasets, these findings motivate future work in which inference for the proposed Bayesian hierarchical model is conducted using variational inference, potentially achieving computational performance comparable to, or possibly exceeding, that of BGLSS.
@@ Acknowledgments
We acknowledge the significant contributions of our coauthor, Ronaldo Dias, who sadly passed away after the completion of this work. His insight, dedication, and expertise were essential to the development of this research.
@@
  • The real data sets are available at the Brasil.IO accessible open data portal (Justen and others, 2021) and IBGE data portal;